OpenStudy (anonymous):

transform quadratic function in standard quadratic form. identify a,b,c 1.) f(x)=(x+1)^2+(x-1)^2

OpenStudy (anonymous):

f(x)=2x^2+0x +2 a=2 b=0 c=2

OpenStudy (anonymous):

need solution how did you get that answer?! solution please?

OpenStudy (anonymous):

ok come to twiddla

OpenStudy (anonymous):

http://www.twiddla.com/587877

OpenStudy (anonymous):

just expend the two and combine the similar ones

OpenStudy (anonymous):

solution please .

OpenStudy (anonymous):

come to twiddla plz

OpenStudy (anonymous):

http://www.twiddla.com/587877

OpenStudy (anonymous):

f(x)=(x+1)^2+(x-1)^2=x^2+2x+1+x^2-2x+1=2x^2+2

OpenStudy (anonymous):

but wait !

OpenStudy (anonymous):

formula for expansion (a+b)^2=a^2+2ab+b^2 (a-b)^2=a^2-2ab+b^2

OpenStudy (anonymous):

i think u will be able to solve it by these formulas

OpenStudy (anonymous):

\[f(x)=(x+1)^2+(x-1)^2\] \[f(x)=x^2+2x+1+x^2-2x+1\] \[f(x)=2x^2+2\] tada